A ball P of mass 2 kg undergoes an elastic collision with another ball Q at rest.After collision, ball P continues to move in its original direction with a speed one fourth of its original speed.What is the mass of the ball Q?


1
Expert's answer
2020-07-22T10:33:46-0400

"mv=mv/4+Mu\to2v=v/2+Mu\to Mu=1.5v"


"mv^2/2=m(v/4)^2/2+Mu^2\to mv^2=mv^2/16+M^2u^2/M"


"2v^2=v^2/8+(1.5v)^2/M\to 2=1/8+2.25/M\to M=1.2kg" Answer.